# Improving Research Idea Generation Through Data: An Empirical Investigation in Social Science

### v1

@@ 4.4 Human Evaluation
w. Validation Tie w/o Validation
Significance 37.5 27.5 35.0
Novelty 45.0 21.7 33.3
Feasibility 40.0 33.3 26.7
Exp. Effectiveness 43.3 27.5 29.2
Overall 42.5 21.7 35.8
[…]
@@
As shown in Table 4, ideas selected with validation processes are ranked higher across all dimensions, with the largest improvement observed in feasibility. This aligns with the reference-based evaluation results and suggests that validation processes provide a valuable signal for enhancing idea selection.
@@ 5.2 Results
(a) Human comparison results of ideas proposed by participants with vs. without access to references (including LLM-generated ideas, related data and validation processes).
Helpfulness High Medium Not Helpful
@@
6 Related Work
Research Idea Generation
There is growing interest in leveraging LLMs for research idea generation, either as a standalone task (Si et al., 2024; Baek et al., 2024) or as part of an end-to-end automated research pipeline (Li et al., 2024b; Lu et al., 2024; Jansen et al., 2025). The former line of work focuses on enhancing the literature search and idea generation stages, typically generating ideas grounded in prior literature (Wang et al., 2024; Yang et al., 2024a). The latter line of work proposes more comprehensive frameworks that extend to later stages, such as experiment design, execution, and paper writing.
Our work aligns more closely with the first line, but novelly incorporates data into the idea generation process. While our work also involves code generation and execution for hypothesis validation, this is not intended as rigorous experimental verification, but serves as a preliminary signal to support idea selection.
Hypothesis Generation
A related but distinct line of work focuses on hypothesis generation, where models generate hypotheses to explain phenomena given access to data, like inducing rules from observations (Zhong et al., 2023; Qiu et al., 2024). Studies in this area explore data-driven methods (Majumder et al., 2024a; Zhou et al., 2024) or integrate literature with data (Liu et al., 2024a), but their goal is to uncover patterns in existing datasets, contrasting with our objective of generating high-quality research ideas.
7 Conclusion and Discussion
Our study shows that incorporating data into research idea generation, through metadata and automatic validation, improves the overall quality of research ideas generated by LLMs. By guiding idea generation with dataset descriptions and selecting ideas given automatic validation processes, LLMs are able to propose ideas that are more feasible and more likely to be effective. Beyond quality improvements, we find that these LLM-generated ideas, along with their validation traces, can serve as valuable inspiration for human researchers.
Limitations
Task Scope
Exploration of LLMs and Validation Methods
Trade-off between Novelty and Feasibility
Ethical Considerations
@@ Appendix A Data Collection
Natural resources rent Rents from coal, oil and natural gas production (of GDP).
@@
1 The Multifaceted Nature of Global Climate Change Negotiations (Bagozzi, 2015)
2 A Closer Look at the Information Provision Rationale: Civil Society Participation in States’ Delegations at the UNFCCC (Böhmelt, 2013)
3 Sectors, Pollution, and Trade: How Industrial Interests Shape Domestic Positions on Global Climate Agreements (Genovese, 2019)
4 The domestic politics of international climate commitments: which factors explain cross-country variation in NDC ambition? (Tørstad et al., 2020)
5 Which Countries Send More Delegates to Climate Change Conferences? Analysis of UNFCCC COPs, 1995–2015 (Kaya and Schofield, 2020)
6 The Institutionalization of a Cleavage: How Differential Treatment Affects State Behavior in the Climate Negotiations (Castro and Kammerer, 2021)
7 How Do Countries Frame Climate Change? A Global Comparison of Adaptation and Mitigation in UNFCCC National Communications (Wright et al., 2023)
8 Institutional Roots of International Alliances: Party Groupings and Position Similarity at Global Climate Negotiations (Genovese et al., 2023)
@@ 1 The Multifaceted Nature of Global Climate Change Negotiations (Bagozzi, 2015)
National communications, highlevel statements, and business statements are collected from the UNFCCC websitehttps://unfccc.int/, which allows free download and copy. Earth negotiation bulletins are collected from the ENB website§§§https://enb.iisd.org/. Meeting attendance records are from Blinova et al. (2024) under the CC0 1.0 license. Democracy index is from Marshall et al. (2014). World risk index is from Welle and Birkmann (2015) under the CC BY license. ND-GAIN vulnerability index is from Chen et al. (2015). All other data in ClimateDataBank are from World Bank Open Data under the CC-BY 4.0 license.
@@ Appendix B Research Topics
Table 8: Climate negotiation research topics used in this paper.
@@ Appendix C Research Idea Generation Methods
  • AI-Researcher (Si et al., 2024): This method first retrieves papers related to the given research topic from Semantic Scholar, uses the retrieved papers to ground idea generation, produces a large number of candidate ideas, and then ranks them to identify the best ones.
  • GPT-Researcher (Elovic, 2023): This method builds a multi-agent framework consisting of planner, executor, and publisher agents. The planner generates plans, while the executor gathers relevant information. The publisher aggregates all information and generates the research ideas.
  • Chain-of-Ideas (Li et al., 2024a): This method enhances the literature search module by organizing relevant literature in a chain structure to effectively mirror the progressive research development.
@@ Appendix E Implementation Details
Appendix F Experimental Results
Gemini-1.5-Pro as the Judge
AI-Researcher 902 933 1047 951 958
938 931 1098 997 991
GPT-Researcher 1019 1000 1045 1015 1020
1073 1021 1183 1134 1103
Chain-of-Ideas 974 1025 822 915 934
1094 1091 805 988 995
Claude-3.5-Sonnet as the Judge
AI-Researcher 870 968 1060 881 945
855 859 1152 918 946
GPT-Researcher 931 972 1076 928 977
1000 903 1228 1085 1054
Chain-of-Ideas 1066 1118 768 1012 991
1278 1180 716 1176 1088
Table 10: ELO scores for research idea generation methods with and without metadata. Better results of each method are in bold.
Table 10 presents the ELO scores for research idea generation methods with and without metadata, serving as the tabular counterpart to Figure 3.
@@ Appendix G Case Study
Table 12: Examples of the automatic validation process.
@@ 2. Output Requirements:
Tables 15 through 17 display the prompts for the automatic validation process, including feasibility checks, hypothesis validation, and validation process summarization. Table 18 outlines the prompt for idea selection in §4, which differs from Table 14 by incorporating the validation process.
Appendix I Annotation Details
@@
Figures 5 and 6 show the annotation interfaces for human evaluations of idea pairs (in §3 - §5) and hypothesis validation processes (in §4.1), respectively. All annotators are fairly paid with more than \$10 per hour.

### v2

@@ 5.3 Human Evaluation
Table 5: Human evaluation of ideas selected by Claude-3.5-Sonnet with and without validation processes ($\%$). P-values are computed with two-tailed Z-tests (* for $p<0.05$, and ** for $p<0.01$).
w. Validation Tie w/o Validation P-value
Significance 37.5 27.5 35.0 0.69
Novelty 45.0 21.7 33.3 0.06
Feasibility 40.0 33.3 26.7 0.03*
Exp. Effectiveness 43.3 27.5 29.2 0.02*
Overall 42.5 21.7 35.8 0.29
@@
As shown in Table 5, ideas selected with validation processes are ranked higher across all dimensions, with the largest improvements observed in feasibility and expected effectiveness. This aligns with the reference-based evaluation results and suggests that validation processes provide a valuable signal for enhancing idea selection.
@@ 6.2 Results
Table 6: Human study results on the inspirational value of LLM-generated ideas ($\%$).
(a) Human comparison of ideas proposed by participants with vs. without access to references (including LLM-generated ideas, related data and validation processes).
@@
High Medium Not Helpful
@@
7 Conclusion
Facing the challenge of generating feasible and effective research ideas with LLMs, we propose a framework that incorporates data into research ideation through metadata and automatic validation. Experiments show that by guiding idea generation with dataset descriptions and selecting ideas given automatic validation processes, LLMs are able to propose ideas that are more feasible and more likely to be effective. Beyond quality improvements, we find that these LLM-generated ideas, along with their validation traces, can serve as valuable inspiration for human researchers.
@@ Appendix A Data Collection
Natural resources rent Rents from coal, oil and natural gas production (% of GDP).
@@
1 The Multifaceted Nature of Global Climate Change Negotiations [3]
2 A Closer Look at the Information Provision Rationale: Civil Society Participation in States’ Delegations at the UNFCCC [5]
3 Sectors, Pollution, and Trade: How Industrial Interests Shape Domestic Positions on Global Climate Agreements [10]
4 The domestic politics of international climate commitments: which factors explain cross-country variation in NDC ambition? [28]
5 Which Countries Send More Delegates to Climate Change Conferences? Analysis of UNFCCC COPs, 1995–2015 [13]
6 The Institutionalization of a Cleavage: How Differential Treatment Affects State Behavior in the Climate Negotiations [6]
7 How Do Countries Frame Climate Change? A Global Comparison of Adaptation and Mitigation in UNFCCC National Communications [31]
8 Institutional Roots of International Alliances: Party Groupings and Position Similarity at Global Climate Negotiations [9]
@@ 1 The Multifaceted Nature of Global Climate Change Negotiations [3]
National communications, highlevel statements, and business statements are collected from the UNFCCC website [^3] https://unfccc.int/, which allows free download and copy. Earth negotiation bulletins are collected from the ENB website [^4] https://enb.iisd.org/. Meeting attendance records are from Blinova et al. [4] under the CC0 1.0 license. Democracy index is from Marshall et al. [23]. World risk index is from Welle and Birkmann [30] under the CC BY license. ND-GAIN vulnerability index is from Chen et al. [7]. All other data in ClimateDataBank are from World Bank Open Data under the CC-BY 4.0 license.
@@ Appendix B Research Topics
Table 9: Climate negotiation research topics used in this paper.
@@ Appendix C Research Idea Generation Methods
  • AI-Researcher [26]: This method first retrieves papers related to the given research topic from Semantic Scholar, uses the retrieved papers to ground idea generation, produces a large number of candidate ideas, and then ranks them to identify the best ones.
  • GPT-Researcher [8]: This method builds a multi-agent framework consisting of planner, executor, and publisher agents. The planner generates plans, while the executor gathers relevant information. The publisher aggregates all information and generates the research ideas.
  • Chain-of-Ideas [15]: This method enhances the literature search module by organizing relevant literature in a chain structure to effectively mirror the progressive research development.
@@ Appendix E Implementation Details
Appendix F Pilot Study: Can LLMs conduct preliminary hypothesis validation?
(a) Accuracy of the hypothesis validation results.
Domain # Papers # Hypotheses Accuracy ($\%$)
Diverse 20 100 78.0
Climate 8 18 72.2
(b) Human evaluation of the validation processes.
Choice Ratio ($\%$)
Mostly Validate the Hypothesis 50
Partially Validate the Hypothesis 40
Does not Help Validating 10
(c) Error analysis of the validation processes. Numbers indicate the ratio of validation processes that encounter the error ($\%$).
30.0 63.3 30.0 33.3
Experimental Setup
We extract 18 hypotheses from the 8 domain-specific papers collected in §3, of which 10 are supported and 8 are refuted. Hypotheses with insignificant or mixed evidence are excluded. To expand the experimental scope, we sample 50 hypotheses from DiscoveryBench [22], drawn from 20 papers across diverse fields like humanities, sociology, and economics. Since all these hypotheses are supported in the original papers, we create 50 negative hypotheses by modifying their variables or relations to balance the evaluation dataset.
We experiment with GPT-4o using the code interpreter assistant, a built-in tool available in GPT models. We input the hypotheses along with their corresponding data into the model. The model engages in multi-turn interactions to write and run Python code in a sandbox environment, to validate whether the hypotheses are supported.
Results
@@ Appendix G Case Study
Table 14: Example of the bottom-ranked idea under the topic Role of Non-State Actors, comparing to the top-ranked idea shown in Table 13.
Topic Role of Non-State Actors
Idea with the Lowest Ranking
Research Question How do non-state actors contribute to the perceived legitimacy of international climate agreements, and what factors enhance their effectiveness in this role?
Theory The legitimacy of climate agreements depends not only on state actors but also on the active participation of NSAs, which represent diverse stakeholder interests. NSAs enhance legitimacy by advocating for justice, inclusivity, and transparency in climate negotiations. Their effectiveness in this role is influenced by their organizational capacity, networks, and alignment with global norms.
Hypotheses 1. NSAs with higher levels of organizational capacity (e.g., funding, staff) are more effective in enhancing the legitimacy of climate agreements.
2. NSAs that engage in transnational networks are more likely to influence public perceptions of climate agreements.
3. The inclusion of NSAs in formal negotiation processes positively correlates with higher public support for climate agreements.
Ideas generated with and without metadata
@@
Top-ranked and bottom-ranked ideas selected with access to the automatic validation process
In the idea selection stage, the idea in Table 13 (referred to as Idea A) is ranked the first among ideas related to Role of Non-State Actors. And Table 14 shows the bottom-ranked idea (referred to as Idea B) under the same topic.
While both ideas offer highly significant research avenues, the ability to empirically measure “NDC ambition” and “NSA participation in UNFCCC meetings” makes Idea A’s hypotheses more directly testable. In contrast, Idea B’s focus on “perceived legitimacy” and “public support” for international agreements presents greater methodological challenges. Measuring these concepts often requires extensive surveys or complex qualitative methods, which can introduce variability and impact the rigor of the findings.
@@ 2. Output Requirements:
Tables 17 through 19 display the prompts for the automatic validation process, including feasibility checks, hypothesis validation, and validation process summarization. Table 20 outlines the prompt for idea selection in §5, which differs from Table 16 by incorporating the validation process.
Appendix I Evaluation Details
@@
I.1 Agreement Analysis
We analyze the agreement between human evaluators and automatic judge models in evaluating LLM-generated ideas.
For human annotators, we measure inter-rater agreement using Krippendorff’s alpha, given that annotator assignments vary across idea pairs. The average Krippendorff’s alpha obtained is 0.30, which is much better than a random rating but far from perfect agreement. This moderate agreement reflects the inherent subjectivity in assessing the quality of research ideas, which is also observed in prior research.
Regarding the two LLM judges in §4, their average Krippendorff’s alpha is 0.24, slightly lower than that observed for human evaluators. However, when we aggregate their individual predictions and computed ELO scores, the Pearson correlation between their generated ELO scores is 0.67. This demonstrates a moderate to high correlation between the two LLM judges’ overall assessments, suggesting a consistent ranking capability despite some individual disagreement on specific idea pairs.
I.2 Annotation Interfaces
Figures 2 and 3 show the annotation interfaces for human evaluations of idea pairs (in §4 - §6) and hypothesis validation processes (in §F), respectively. All annotators are fairly paid with more than \$10 per hour.
@@ I.2 Annotation Interfaces
Appendix J Limitations
Task Scope
Exploration of LLMs and Validation Methods
Trade-off between Novelty and Feasibility
The Role of LLM Ideation
Appendix K Ethical Considerations